if you had 5 pounds of something and it coast 2.79 how much will it coast to have 50 pounds ? may you please help me solve this question

Answer by TimothyLamb(4379)

cpp = 2.79/5
---
2.79/5 = c/50
c = (2.79/5) * 50 = $27.90
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php